OpenStudy (anonymous):

one positive integer is 4 less than another. the product of the two integers is 12. what are the integers?

OpenStudy (mathstudent55):

Are you learning systems of equations?

OpenStudy (anonymous):

Solving Problems With Quadratic Equations.. i need major help..

OpenStudy (mathstudent55):

Ok. I'll help you. We need to write an equation first. You have 2 numbers, one is 4 less than the other.

OpenStudy (mathstudent55):

Let x be the smaller of the two numbers. What is the larger number in terms of x? Remember that if the smaller number is 4 less than the larger number, then the larger number is 4 more than the smaller number.

OpenStudy (anonymous):

i honestly do not know any of this.. im struggling so bad..

OpenStudy (mathstudent55):

Ok. Let's say a number is 6. Then a number that is 4 more than 6 is 6 + 4 (which happens to be 10). In our case, we don't know what the numbers are. We are calling the smaller number x since it is unknown to us. If the smaller number is x, then what is a number that is 4 more than x? Do the same I did with 6 and 4 above.

OpenStudy (mathstudent55):

What operation do you do to any number to end up with a number that is 4 more than the first number?

OpenStudy (anonymous):

addition

OpenStudy (anonymous):

are you there?..

OpenStudy (mathstudent55):

Yes. It's addition. If you add 4 to any number, you end up with a number that 4 more than the first number. Here, we have chosen x to be the smaller number. Then 4 more than x is the larger number. 4 more than x is x + 4 That means the two numbers are x and x + 4 Do you follow so far?

OpenStudy (mathstudent55):

Are you still there?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Do you understand that the numbers are x and x + 4?

OpenStudy (anonymous):

yes i understand how you got them

OpenStudy (mathstudent55):

Ok. Now we follow the rest of the problem. The problem states " the product of the two integers is 12." Since the integers are x and x + 4, we multiply them and set the product equal to 12. x(x + 4) = 12 Ok so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you there

OpenStudy (mathstudent55):

My connection is messed up. I keep losing it.

OpenStudy (mathstudent55):

Now you take the equation we have, multiply out the left side, subtract the number from the right side to the left, and you'll end up with a quadratic equation.

OpenStudy (mathstudent55):

x(x + 4) = 12 We use the distributive property on the left side to get this: \(x^2 + 4x = 12\) Now we subtract 12 from both sides: \(x^2 + 4x - 12 = 0\) Now we need to solve the quadratic equation. You can try factoring. You need two numbers that add to 4 and multiply to -12.

OpenStudy (anonymous):

2,-6??

OpenStudy (anonymous):

oh wow really?

OpenStudy (mathstudent55):

2 and -6 do multiply to -12, but they add to -4. You need two numbers that multiply to -12 and add to 4, not -4. You're close. If you just change the signs of your two numbers, you'll get them.